OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1 = 500, an = (an-1)/5

OpenStudy (jtvatsim):

We are given \[a_1 = 500, \ \ a_n = \frac{a_{n-1}}{5}\] and asked to find the first five terms. Is that the question?

OpenStudy (anonymous):

yea

OpenStudy (jtvatsim):

OK, this one is sort of easy and fun once you understand what all the formulas mean. :)

OpenStudy (anonymous):

Im ready to take notes lol

OpenStudy (jtvatsim):

Mathematicians use the \[a_n\] as a way to stand for the nth term. That is, \[a_1\] is the 1st term, \[a_2\] is the 2nd term, \[a_3\] is the 3rd term, and so on...

OpenStudy (jtvatsim):

Does that make sense so far?

OpenStudy (anonymous):

yes so i replace n with 1, 2, 3 etc?

OpenStudy (jtvatsim):

Exactly! And you were already given the first term, that is, \[a_1 = 500\]

OpenStudy (jtvatsim):

The other formula tells you how to find the next term.

OpenStudy (jtvatsim):

So, \[a_n = \frac{a_{n-1}}{5}\] means to find the next term "an", take the previous term "an-1" and divide by 5

OpenStudy (jtvatsim):

So, what happens when you plug in n = 2?

OpenStudy (anonymous):

soo its 500-1/5 ?

OpenStudy (jtvatsim):

Almost. :) Here's the way it works, you are close...

OpenStudy (jtvatsim):

\[a_n = \frac{a_{n-1}}{5}\] when I plug in n = 2, I see \[a_2 = \frac{a_{2-1}}{5}\] which is \[a_2 = \frac{a_1}{5}\]

OpenStudy (anonymous):

whats a ?

OpenStudy (jtvatsim):

a isn't really anything by itself... a1 is the first term, a2 is the second term, a3 is the third term and so on, like we looked at before.

OpenStudy (jtvatsim):

What the equation says is that a2 is equal to a1 divided by 5. Do you see that?

OpenStudy (anonymous):

yea i see

OpenStudy (jtvatsim):

In english, the second term is equal to the first term divided by 5.

OpenStudy (anonymous):

but where does the 500 go ?

OpenStudy (jtvatsim):

Good question!

OpenStudy (jtvatsim):

So, since \[a_1 = 500\], we just get \[a_2 = \frac{a_1}{5}=\frac{500}{5}=100\]

OpenStudy (jtvatsim):

See how it works?

OpenStudy (anonymous):

kinda

OpenStudy (jtvatsim):

It is a bit tricky at first, but you'll get it soon. :)

OpenStudy (anonymous):

no then on the next one would be 100/5?

OpenStudy (anonymous):

so*

OpenStudy (jtvatsim):

Yes, that is right. The third term is the second term divided by 5 (according to the formula)

OpenStudy (jtvatsim):

In math lingo, \[a_3 = \frac{a_2}{5}\]

OpenStudy (jtvatsim):

So since \[a_2 = 100\] we have to have \[a_3 = \frac{100}{5} = 20\]

OpenStudy (anonymous):

okay heres what i have so far 100, 20, 4, 4/5?

OpenStudy (jtvatsim):

Those are the 2nd, 3rd, 4th, and 5th terms, good that is accurate.

OpenStudy (anonymous):

and then 4/25?

OpenStudy (jtvatsim):

That would be the 6th term, yes.

OpenStudy (jtvatsim):

But the question only wants us to find the first 5 terms.

OpenStudy (jtvatsim):

So, we just need the list for the 1st, 2nd, 3rd, 4th, and 5th terms.

OpenStudy (anonymous):

oooh ! i get it ! 500 is the first

OpenStudy (jtvatsim):

YES!!! :D

OpenStudy (anonymous):

yay omgee i was putting the wrong answer :p

OpenStudy (anonymous):

thank you so much (:

OpenStudy (jtvatsim):

Oh, haha, I've done that before too. ;) no problem!

OpenStudy (jtvatsim):

Good job, you are better at math than you think. :)

OpenStudy (anonymous):

haha i just people to explain it to me :p cause i get so lost sometimes

OpenStudy (jtvatsim):

You aren't the first to get lost... math class is made a lot harder than it needs to be... sometimes plain english is better than "math speak" :)

OpenStudy (anonymous):

haha it is :p

OpenStudy (anonymous):

thank you again

OpenStudy (jtvatsim):

no problem, good night! :)